Use FOIL to multiply:
(5x – 1)(3x – 2)

Answer by holly9010(3)

F= 15x
O= -10x
I= -3x
L= 2
15x-10x-3x+2=0
5x-3x+2=0
2x+2=0
2x=-2
x=-1